OpenStudy (hulkart):

Given the expression (1 + 2i)2, perform the indicated operation and write the answer in the form a + bi. −1 − 2i −3 + 4i −4 − 4i −3 − 4i

OpenStudy (hulkart):

yes they are.

OpenStudy (hulkart):

@iGreen help pls

OpenStudy (welshfella):

(1 + 2i)(1 + 2i) = (1 + 2i) + 2i(1 + 2i) can you continue?

OpenStudy (welshfella):

distribute 2i over (1 + 2i)

OpenStudy (welshfella):

= 1 + 21 + 2i*1 + 2i*2i = ?

OpenStudy (welshfella):

8 sorry that's 1 + 2i + 2i*1 + 2i*2i

OpenStudy (welshfella):

what is 2i* 2i?

OpenStudy (welshfella):

no its 4i^2 do you know what the definition of i is?

OpenStudy (welshfella):

when you manipulate it algebraically just treat it as a variable like x or y.

OpenStudy (welshfella):

its an imaginary number and equal to sqrt(-1) so i^2 = sqrt(-1) squared = -4 so 4i^2 = -4

OpenStudy (welshfella):

gotta go right now

OpenStudy (radar):